# Spatiotemporal-Augmented Graph Neural Networks for Human Mobility Simulation

### v1

@@ 5.1 Datasets
To ensure reproducibility and facilitate fair comparisons with previous works, we evaluates the human mobility of four cities (i.e., New York, Tokyo, Moscow and Singapore) extracted from the publicly available dataset Foursquare in line with previous studies [14, 64]. The timespan and the numbers of users, locations and visit records in each dataset are shown in Table I. The wide range of the timespan and the large-scale visits can sufficiently compare the STAR framework against baseline methods. In experiments, we use the original raw datasets that only contain the GPS coordinates of each location and user check-in records, and pre-process them following the protocol of the human mobility simulation task.
Specifically, we split the whole dataset into three parts: a training set for training the generative model, a validation set for finding the best parameters of models and a testing set for the final evaluation of various metrics. The partition of the four datasets is all set as 7:1:2. Besides, we set the basic time slot as an hour of the day for the convenience and universality of modeling. Finally, in order to ensure the effectiveness and accuracy of modeling, we only remain the trajectories with visit records greater than eight each day.
@@ 5.2 Baseline Methods and Evaluation Metrics
  • DeepMove [12] learns the periodic patterns of mobility by the recurrent network and the neural attention layer.
@@
  • SeqGAN [33] solves the discrete sequence generation problem based on the policy gradient technique with GAN, which is suitable for trajectory generation.
@@ 5.3 Experimental Settings
For a fair comparison, we set all methods with the following settings: the hidden dimension of embeddings is set to 32, the number of epochs is set to 50, and the batch size is set to 32. Other specific hyper-parameters of these baselines follow the best settings in their papers. The proposed STAR framework is implemented in PyTorch [67]. In default settings, the learning rate is set to 0.01 and the dropout ratio is set to 0.6. The number of attention layers is searched over {1, 2}, and the number of attention heads is searched over {1, 2, 4}.
@@ 5.4 RQ1: Performance Comparison
  • Our framework steadily achieves the best performance. STAR achieves the best performance on all datasets with ranking first on nineteen metrics and ranking second on two metrics over twenty four metrics of the four datasets. For the nineteen ranking first metrics, compared with the best baseline, our method reduces the JSD up to 53.50%. For the other five metrics, namely Duration on the NYC dataset, DailyLoc on the Moscow dataset, I-rank on the NYC, TKY and Singapore dataset, our framework also obtains competitive performance with the best baseline.
  • Model-based methods are limited in simulating human mobility. Markov performs worse on the time-dependent metrics (i.e., duration and dailyloc) but better on the distance-based metrics (i.e., Distance and Radius), because it obtains the next location based on the distribution of historical transition probabilities, which is in line with our intuition that individuals are prone to the near areas. The performance of IO-HMM is also unsatisfactory, because its modeling relies on extensive manual labeling, which places higher demands on the quality of data. For example, lack of records at home interferes with the annotation of Home label, which degrades the predictive performance of IO-HMM. In addition, the sparsity of data introduces errors in the labeling of dwell time.
@@
  • Generative networks fail to generate realistic human trajectories without trajectory pre-training. GAN performs the worst across almost all metrics, indicating that it is difficult to capture the hidden patterns of human mobility when learning with noisy and inaccurate raw data. SeqGAN is pre-trained by the task based on human mobility modeling, so it yields much better results than GAN.
  • It is essential to model dynamic spatiotemporal dependencies among locations. Although CGE leverages graph structure for data augmentation, its graph is static and the node embeddings generated by Word2vec cannot be learned and updated in the process of mobility simulation, so it only learns better on several metrics but performs poorly overall. MoveSim introduces the urban structure modeling component especially for locations, so it achieves the best result on Duration metric of the NYC dataset and the second-best result on I-rank metric of the TKY and Moscow dataset, radius metric of the Moscow dataset and Duration metric of the Singapore dataset.
@@ 5.5 RQ2: Ablation Study of STAR
Figure 3: Ablation study on the channels of graphs. STG, SDG and TTG represent SpatioTemporal Graph, Spatial Distance Graph and Temporal Transition Graph respectively. MSC and SGP are short for Moscow and Singapore. A lower JSD value indicates a better performance.
TABLE III: Ablation study on edge type of GNNs in STAR in terms of JSD. MSC and SGP are short for Moscow and Singapore. A lower JSD value indicates a better performance.
@@
NYC Weighted 0.1117 0.3665 0.0013 0.0613 0.0383 0.0628
Vanilla 0.1105 0.3636 0.0011 0.0580 0.0359 0.0624
% Improv. 1.07% 0.79% 15.38% 5.38% 6.27% 0.64%
TKY Weighted 0.1247 0.4220 0.0006 0.0438 0.0298 0.0697
Vanilla 0.1196 0.4178 0.0005 0.0402 0.0259 0.0693
% Improv. 4.09% 1.00% 16.67% 8.22% 13.09% 0.57%
MSC Weighted 0.0081 0.0201 0.0029 0.0501 0.0335 0.0720
Vanilla 0.0049 0.0152 0.0023 0.0497 0.0294 0.0485
% Improv. 39.51% 24.38% 20.69% 0.80% 12.24% 32.64%
SGP Weighted 0.0098 0.0115 0.0012 0.0384 0.0295 0.0494
Vanilla 0.0062 0.0095 0.0014 0.0371 0.0287 0.0485
% Improv. 36.73% 17.39% -16.67% 3.39% 2.71% 1.82%
TABLE IV: Ablation study on dwell branch in STAR in terms of JSD. DB, MSC and SGP are short for the dwell branch, Moscow and Singapore. A lower JSD value indicates a better performance.
@@
TKY w/o DB 0.1237 0.4293 0.0004 0.0433 0.0343 0.0705
STAR 0.1196 0.4178 0.0005 0.0402 0.0259 0.0693
% Improv. 3.31% 2.68% -25.00% 7.16% 24.49% 1.70%
MSC w/o DB 0.0076 0.0180 0.0027 0.0501 0.0297 0.0555
STAR 0.0049 0.0152 0.0023 0.0497 0.0294 0.0485
% Improv. 35.53% 15.56% 14.81% 0.80% 1.01% 12.61%
SGP w/o DB 0.0065 0.0101 0.0016 0.0384 0.0320 0.0497
STAR 0.0062 0.0095 0.0014 0.0371 0.0287 0.0485
% Improv. 4.62% 5.94% 12.50% 3.39% 10.31% 2.41%
Figure 4: Effects of the number of layers and attention heads in STAR.
@@ 5.5.1 Designs of Multi-channel Graph
In order to learn the spatiotemporal correspondence among locations, we design three kinds of spatiotemporal graphs in the multi-channel embedding module to represent the node embeddings separately and obtain the final embeddings with a fusion layer. To verify the effectiveness of the three graphs in STAR, we get rid of any of them (i.e., without SpatioTemporal Graph, without Spatial Distance Graph and without Temporal Transition Graph) to learn the node embeddings and perform human mobility simulation. The comparison of their simulated results is shown in Figure 3.
Compared with the results predicted by removing any of the three graphs, the fused embedding achieves optimal performance on almost all metrics over the four datasets. In addition, the average performance of removing SpatioTemporal Graph is the worst on six metrics, especially on Distance, Duration, DailyLoc and G-rank metric, which measure the effectiveness of simulated trajectories from both spatial and temporal perspectives. It indicates that only relying on the static distance and transition information of locations is not enough to carry out accurate trajectory simulation.
Eliminating the Spatial Distance Graph results in a decline of performance on distance-based metrics (i.e. Distance and Radius), which is as per our hypothesis. However, it scores well on preference-based metrics (i.e. G-rank and I-rank), which is attributed to the SpatioTemporal Graph and Temporal Transition Graph effectively learning frequently visited locations in the trajectories. Furthermore, removing the Temporal Transition Graph results in inferior performance compared to remaining all three graphs but surpasses removing any other graph overall, which suggests that the SpatioTemporal Graph can aptly supplement the temporal transition information of trajectories.
@@ 5.5.2 Designs of Edge Type in GNNs
In order to explore the effect of edge weight in GNNs on model performance, we set two edge types for model, Weighted and Vanilla. Weighted retains the weights of edges in the three graphs, that is, the three graphs before binaryzation, while Vanilla only uses the adjacency relationship of the three graphs to learn node embedding. We list their performance in Table III.
It can be seen from the results that Vanilla is better than Weighted on almost all metrics of the four datasets, especially on the Distance metric of the Moscow and Singapore dataset, the improvement reaches 39.51% and 36.73%, respectively. The reasons for the result may be as follows. Firstly, Vanilla are simpler to implement, while Weighted increases the training time and degrades the performance. Secondly, Weighted edges may assign significant weights to noisy or irrelevant features, leading to reduced model performance. In contrast, Vanilla only consider the presence or absence of an edge, which can be more resilient to noise.
The six metrics can be classified into three categories based on the level of improvement they show when comparing the Vanilla graph to the Weighted graph. Distance-based metrics (i.e. Distance and Radius) exhibit the highest improvement, whereas time-dependent metrics (i.e. duration and dailyloc) show the least improvement. Preference-based metrics (i.e. G-rank and I-rank) show moderate improvements. The underlying reason lies in the fact that our proposed multi-channel graphs explicitly encode the distance and user preferences of locations, while lack the time-dependent constraints. Additionally, the long-tailed distribution of the Weighted version might exaggerate the uneven distribution of location frequencies.
@@ 5.5.3 Designs of Dwell Branch
The dwell branch which allows individuals to remain in the current location with a learnable probability, aims to adaptively perceive the varying durations in mobility trajectories. We conduct ablation experiments on this component to verify its effectiveness, and the results are shown in Table IV.
As we can observe, the design of the dwell branch enhances the performance on twenty three metrics over twenty four metrics of the four datasets, with the most significant improvement observed with the Distance metric on the Moscow dataset, which is up to 35.53%. It implies that enabling individuals to dwell in a specific location with a learnable probability can facilitate a more comprehensive representation of complex real-world behavior, consequently improving the performance of various metrics from different perspectives.
Specifically, there are two reasons for the performance improvement. First of all, staying at the current location with a certain probability reduces the likelihood of long-distance movement for a individual. As a result, the dwell branch facilitates the acquisition of distance-related patterns and features, subsequently improving the performance of distance-based and time-dependent metrics. Secondly, individuals tend to visit several specific locations frequently and thereby the learnable stay probability can increase individuals’ visit probability at frequently visited places, which effectively captures the mobility preferences of real trajectories from both collective and individual view.
It is worth noting that the dwell branch achieves the most significant performance improvement on the Moscow dataset, as expected. Due to the high frequency of staying in the previous location in the Moscow dataset, the learnable stay probability is essential for modelling repeated movement patterns within the trajectories more effectively.
@@ 5.6 RQ3: Parameter Sensitivity of STAR
Parameter sensitivity analysis can help us understand how changes in parameters of model affect its performance, which is conducive to optimizing the model for better performance, identifying key parameters that have the most important impact on results, and assessing the reliability and robustness of the model. Figure 4 demonstrates the parameter sensitivity of STAR on two key hyper-parameters, the number of GAT layers and attention heads, which affect the graph convolution architecture in the STAR framework.
We plot the grid search results of the number of GAT layers over {1, 2} and the number of attention heads over {1, 2, 4} to thoroughly test the STAR framework on twenty four metrics over the four datasets. As can be seen from Figure 4, the performance of STAR is robust under different hyperparameter settings, which indicates different hyperparameter values wouldn’t affect its superiority over other baselines.
Specifically, the performance of different hyperparameters varies from metric to metric. On distance-based metrics (i.e., Distance and Radius), only a simple model with a few number of GAT layers and attention heads can achieve good performance. As the number of layers and attention heads increases, the performance of the model on distance-based metrics deteriorates due to overfitting. The phenomenon emerges since moving distance is the most conspicuous pattern of trajectories, where the distance-related features in trajectories can be effectively modeled with fewer GAT layers and attention heads. On time-dependent metrics (i.e., duration and dailyloc), the performance of the model exhibits minimal fluctuations with the changes in the number of GAT layers and attention heads, which demonstrates the robustness of our proposed method.
On preference-based metrics (i.e., G-rank and I-rank), a simple model with a few number of GAT layers and attention heads performs poorly. On the one hand, deeper GAT layers improves the performance of the model, which is because deeper layers can effectively capture complex and advanced graph structures. By using multiple GAT layers, the model can handle increasingly abstract representations of the input graphs, thereby allowing the model to learn more advanced features and make better simulation. In addition, deeper GAT layers can improve the model’s ability to learn relationships between remote nodes, which is important for capturing dynamic spatio-temporal dependencies among locations. On the other hand, the increase in the number of attention heads also improves the model’s performance. The reason for the result is that having more attention heads improves the quality of the learned node representation. Each attention head learns a different attention weight, enabling the model to capture different and potentially complementary information from the neighborhood of each node. By integrating information from multiple attention heads, the model can learn a richer representation of the graph and produce better simulation. Furthermore, the increase in the number of attention heads can make the model more resilient to different types of input graphs, improving its adaptability to unknown graph structures.
However, when the number of GAT layers is 2 and the number of attention heads increases to 4, the model performance decreases, which is even worse than the model with one layer and four attention heads. As mentioned above, too many attention heads tend to overfit, where the extra head no longer improves or even degrades the performance of the model.
Figure 5: The geographical visualization of location visit frequency of real and simulated human trajectories on the TKY dataset.

### v3

@@ 5.1 Datasets
To ensure reproducibility and facilitate fair comparisons with previous works, we evaluate the human mobility of four cities (i.e., New York, Tokyo, Moscow and Singapore) extracted from the publicly available dataset Foursquare in line with previous studies [14, 69]. The timespan and the numbers of users, locations and visit records in each dataset are shown in Table I. The wide range of the timespan and the large-scale visits can sufficiently compare the STAR framework against baseline methods. In experiments, we use the original raw datasets that only contain the GPS coordinates of each location and user check-in records, and pre-process them following the protocol of the human mobility simulation task.
Specifically, we split the whole dataset into three parts: a training set for training the generative model, a validation set for finding the best parameters of models and a testing set for the final evaluation of various metrics. The partition of the four datasets is set as 7:1:2. Besides, we set the basic time slot as an hour of the day for the convenience and universality of modeling. Finally, in order to ensure the effectiveness and accuracy of modeling, we only remain the trajectories with visit records greater than eight each day.
@@ 5.2 Baseline Methods and Evaluation Metrics
  • DeepMove [12] employs an attentional recurrent network to acquire the knowledge of sequential transitions and periodic movement patterns from lengthy and sparse trajectories.
@@
  • SeqGAN [35] solves the problem of generating discrete sequence by leveraging the policy gradient technique in conjunction with GAN.
@@
  • VOLUNTEER [68] proposes a user VAE and a trajectory VAE to capture mobility patterns from both group and individual views.
@@ 5.3 Experimental Settings
(1) Baseline methods with adversarial learning process (i.e. GAN, SeqGAN, MoveSim) use a hidden dimension of 32 for generator and a hidden dimension of 64 for discriminator. In addition, they are pre-trained with 100 epochs and adversarially trained with 50 epochs. (2) All other baseline methods use a hidden dimension of 32 and are trained with 50 epochs. (3) All methods use a batch size of 32. The other specific hyperparameters of the baselines follow the settings reported in their respective papers. For STAR, the number of GAT layers is searched over {1, 2, 3}, and the number of attention heads is searched over {1, 2, 4, 8}.
@@ 5.4 RQ1: Performance Comparison
  • Our framework steadily achieves the best performance. STAR achieves the best performance on all datasets with ranking first on twenty-one metrics and ranking second on one metric over twenty-four metrics of the four datasets. For the twenty-one ranking first metrics, compared with the best baseline, our method reduces the JSD up to 60%. For the other three metrics, namely I-rank on the NYC and TKY dataset, G-rank on the Singapore dataset, our framework also obtains competitive performance with the best baseline.
  • Model-based methods are limited in simulating human mobility. Markov performs worse on the time-dependent metrics (i.e., Duration and Dailyloc) but better on the distance-based metrics (i.e., Distance and Radius), because it obtains the next location based on the distribution of historical transition probabilities, which aligns with our intuition that individuals are more likely to visit nearby areas. The performance of IO-HMM is also unsatisfactory, because its modeling relies on extensive manual labeling, which places higher demands on the quality of data. For example, lack of records at home interferes with the annotation of Home label, which degrades the predictive performance of IO-HMM. In addition, the sparsity in data introduces errors when labeling dwell time.
@@
  • Generative networks fail to generate realistic human trajectories without trajectory pre-training. GAN performs the worst across almost all metrics, indicating that it is difficult to capture the hidden patterns of human mobility when learning with noisy and inaccurate raw data. SeqGAN is pre-trained by the task based on human mobility modeling, so it yields much better results than GAN. MoveSim introduces the urban structure modeling component especially for locations, so it achieves the second-best result on Duration metric of NYC dataset.
  • It is essential to model dynamic spatiotemporal dependencies among locations. Although CGE leverages graph structure for data augmentation, its graph is static and the node embeddings generated by Word2vec cannot be learned and updated in the process of mobility simulation, so it only learns better on several metrics but performs poorly overall. VOLUNTEER achieves the second-best results on the Distance and Radius metrics due to its specific modeling for workplace and residence distribution from the group view.
@@ 5.5 RQ2: Ablation Study of STAR
Figure 3: Ablation study on the channels of graphs. All experimental results are conducted over five trials for a fair comparison. STG, SDG and TTG represent SpatioTemporal Graph, Spatial Distance Graph and Temporal Transition Graph respectively. MSC and SGP are short for Moscow and Singapore. A lower JSD value indicates a better performance.
TABLE III: Ablation study on edge type of GNNs in STAR in terms of JSD. All experimental results are conducted over five trials for a fair comparison. MSC and SGP are short for Moscow and Singapore. A lower JSD value indicates a better performance.
@@
NYC Weighted 0.1150 0.3672 0.0014 0.0613 0.0387 0.0584
Vanilla 0.1134 0.3615 0.0012 0.0597 0.0378 0.0550
% Improv. 1.43% 1.56% 14.29% 2.68% 2.33% 5.89%
TKY Weighted 0.1245 0.4256 0.0005 0.0351 0.0326 0.0693
Vanilla 0.1206 0.4198 0.0004 0.0340 0.0307 0.0650
% Improv. 3.13% 1.36% 24.00% 3.25% 5.89% 6.20%
MSC Weighted 0.0103 0.0234 0.0031 0.0524 0.0206 0.0555
Vanilla 0.0071 0.0165 0.0022 0.0516 0.0196 0.0518
% Improv. 31.07% 29.66% 27.74% 1.60% 4.66% 6.63%
SGP Weighted 0.0112 0.0144 0.0011 0.0322 0.0310 0.0427
Vanilla 0.0080 0.0108 0.0009 0.0330 0.0283 0.0418
% Improv. 28.57% 25.14% 16.36% - 8.58% 2.15%
TABLE IV: Ablation study on dwell branch in STAR in terms of JSD. All experimental results are conducted over five trials for a fair comparison. DB, MSC and SGP are short for the dwell branch, Moscow and Singapore. A lower JSD value indicates a better performance.
@@
TKY w/o DB 0.1322 0.4484 0.0007 0.0379 0.0371 0.0644
STAR 0.1206 0.4198 0.0004 0.0340 0.0307 0.0650
% Improv. 8.77% 6.37% 45.71% 10.40% 17.30% -
MSC w/o DB 0.0155 0.0247 0.0041 0.0531 0.0266 0.0606
STAR 0.0071 0.0165 0.0022 0.0516 0.0196 0.0518
% Improv. 54.19% 33.36% 45.37% 2.90% 26.17% 14.49%
SGP w/o DB 0.0123 0.0126 0.0018 0.0341 0.0439 0.0485
STAR 0.0080 0.0108 0.0009 0.0330 0.0283 0.0418
% Improv. 34.96% 14.44% 48.89% 3.17% 35.44% 13.86%
Figure 4: Effects of the number of layers and attention heads on STAR. All experimental results are conducted over five trials.
@@ 5.5.1 Designs of Multi-channel Graph
In order to learn the spatiotemporal correspondence among locations, we design three kinds of spatiotemporal graphs in the multi-channel embedding module to represent the node embeddings separately and obtain the final embeddings with a fusion layer. To verify the effectiveness of the three graphs in STAR, we get rid of any of them (i.e., without SpatioTemporal Graph, without Spatial Distance Graph and without Temporal Transition Graph) to learn the node embeddings and perform human mobility simulation. Additionally, we perform the experiment without any spatiotemporal graph to validate the effectiveness of the entire Multi-channel Embedding Module. The comparison of their simulated results is shown in Figure 3.
Obviously, the performance significantly deteriorates when spatiotemporal graphs are not used (i.e. w/o ALL), which suggests the necessity of incorporating spatiotemporal graphs for locations. Compared with the results predicted by removing any of the three graphs, the fused embedding achieves optimal performance on almost all metrics over the four datasets. In addition, the average performance of removing SpatioTemporal Graph is the worst on six metrics, especially on Distance, Radius, G-rank and I-rank metrics, which measure the effectiveness of simulated trajectories from both spatial and preference perspectives. It indicates that relying solely on the static distance and transition information of locations is not enough to carry out accurate trajectory simulation.
Eliminating the Spatial Distance Graph results in a decline of performance on distance-based metrics (i.e. Distance and Radius), which is as per our hypothesis. However, it scores well on temporal and preference metrics (i.e. Duration and G-rank), which is attributed to the SpatioTemporal Graph and Temporal Transition Graph effectively learning the temporal patterns and the frequently visited locations in the trajectories. Furthermore, removing the Temporal Transition Graph results in inferior performance compared with remaining all three graphs but surpasses removing any other graph overall, which suggests that the SpatioTemporal Graph can aptly supplement the temporal transition information of trajectories.
@@ 5.5.2 Designs of Edge Type in GNNs
In order to explore the effect of edge weight in GNNs on the model performance, we set two edge types for model, Weighted and Vanilla. Weighted retains the weights of edges in the three graphs, that is, the three graphs before binarization, while Vanilla only uses the adjacency relationship of the three graphs to learn node embedding. We list their performance in Table III.
It can be seen from the results that Vanilla is better than Weighted on almost all metrics of the four datasets, especially on the Distance metric of the Moscow and Singapore dataset, the improvement reaches 31.07% and 28.57% , respectively. The reason for the observed result is that the original graph structure may not be optimal or informative [72, 73]. Firstly, Weighted edges may introduce additional noise or bias during the learning process, leading to suboptimal or inaccurate models. Secondly, Vanilla edges can simplify the learning process and yield a more robust and generalizable representation, thereby improving the performance on the human mobility simulation task.
The six metrics can be classified into three categories based on the level of improvement they show when comparing the Vanilla graph to the Weighted graph. Distance-based metrics (i.e. Distance and Radius) exhibit the highest improvement, whereas preference-based metrics (i.e. G-rank and I-rank) show the least improvement. Time-dependent metrics (i.e. Duration and Dailyloc) showcase moderate improvements. The underlying reason lies in the fact that our proposed multi-channel graphs explicitly encode the distance and temporal transitions among locations, while lacking user preference constraints. Additionally, the long-tailed distribution of the Weighted version might exacerbate the uneven distribution of location frequencies.
@@ 5.5.3 Designs of Dwell Branch
The dwell branch which allows individuals to remain in the current location with a learnable probability, aims to adaptively perceive the varying durations in mobility trajectories. To verify its effectiveness, we conduct the ablation experiments shown in Table IV. As we can observe, the design of the dwell branch enhances the performance on twenty three metrics over twenty four metrics of the four datasets, with the most significant improvement observed with the Distance metric on the Moscow dataset, which is up to 54.19%. It implies that enabling individuals to dwell in a specific location with a learnable probability can facilitate a more comprehensive representation of the complex real-world behavior, consequently improving the performance of various metrics from different perspectives.
Specifically, there are two reasons for the performance improvement. First of all, staying at the current location with a certain probability reduces the likelihood of long-distance movement for an individual. As a result, the dwell branch facilitates the acquisition of distance-related patterns and features, subsequently improving the performance of distance-based and time-dependent metrics. Secondly, individuals tend to frequently visit several specific locations and thereby the learnable stay probability can increase individuals’ visit probability at frequently visited places, which effectively captures the mobility preferences of real trajectories from both collective and individual views. It is worth noting that the dwell branch achieves the most significant performance improvement on the Moscow dataset, as expected. Due to the high frequency of staying in the previous location in the Moscow dataset, the learnable stay probability is essential for modelling repeated movement patterns within the trajectories more effectively.
@@ 5.6 RQ3: Parameter Sensitivity of STAR
Parameter sensitivity analysis can help us understand how changes in parameters of model affect its performance, which is conducive to optimizing the model’s performance, identifying key parameters that have the most important impact on results, and assessing the reliability and robustness of the model. We plot the grid search results of the number of GAT layers over {1, 2, 3} and the number of attention heads over {1, 2, 4, 8} to thoroughly test the STAR framework on twenty four metrics over the four datasets. As can be seen from Figure 4, the performance of STAR is robust under different hyperparameter settings, which indicates different hyperparameter values wouldn’t affect its superiority over other baselines.
Specifically, the performance of different hyperparameters varies from metric to metric. On distance-based metrics (i.e., Distance and Radius), only a simple model with a few number of GAT layers and attention heads can achieve good performance. As the number of layers and attention heads increases, the performance of the model on distance-based metrics deteriorates due to overfitting. The phenomenon emerges since moving distance is the most conspicuous pattern of trajectories, where the distance-related features in trajectories can be effectively modeled with fewer GAT layers and attention heads. On time-dependent metrics (i.e., Duration and Dailyloc), the performance of the model exhibits minimal fluctuations with the changes in the number of GAT layers and attention heads, which demonstrates the robustness of our proposed method. On preference-based metrics (i.e., G-rank and I-rank), a simple model with a few number of GAT layers and attention heads performs poorly. Firstly, deeper GAT layers improves the performance of the model, which is because deeper layers can effectively capture complex and advanced graph structures. Secondly, the increase in the number of attention heads also improves the model’s performance. Multiple attention heads enable the model to capture different and potentially complementary information from the neighborhood of each node, improving the quality of the learned node representation.
Figure 5: The geographical visualization of location visit frequency of real and simulated human trajectories on the NYC dataset.